# Searching for star formation towards the Eos molecular cloud

### v1

@@ 3.1 Color Magnitude Diagram
To analyze the age distribution of stars in the Eos Cloud, we constructed a Color-Magnitude Diagram (CMD) using Gaia DR3 photometric data. The CMD plots the absolute G magnitude ($M_{\textrm{G}}$) against the ($G_{\textrm{BP}}$-$G_{\textrm{RP}}$)0 color index, providing a clear visualization of the stellar population in the region. Figure 2 shows a CMD for our extinction, reddening and parallax-corrected stellar sample. Overlaid are Baraffe et al. (2015) PMS evolutionary tracks and isochrones. Other star forming regions on the edge of the local supernova-driven bubble have typical stellar ages $<3$ Myr (e.g. for Taurus, Lupus, Ophiuchus, Chameleon; Comerón, 2008; Galli et al., 2021; Miret-Roig et al., 2022; Luhman, 2025). Conversely, towards the Eos cloud there are no stars at all in that age range over the $70-150$ pc distance we considered (Eos spans 94-136 pc). There is a population of $<20$ Myr stars, however we will demonstrate in section 3.3 that a population similar to this is found in other similar search volumes at this galactic latitude. The Gaia CMD hence demonstrates convincingly that star formation is not currently actively taking place in the Eos cloud.
@@ 3.3 Adjacent Regions
The absence of clustering effects and the uniformity across all the regions as mentioned above, suggest that possibly the stars in the Eos cloud and its surroundings have been dynamically dispersed over time. Possible mechanisms for this dispersion include Galactic tidal interactions, past dynamical mixing, or dissolution of an older stellar association (Bland-Hawthorn & Gerhard, 2016). Future studies incorporating full three-dimensional velocity data (e.g., using radial velocity measurements from Gaia DR4 or spectroscopic surveys) could provide deeper insights into the origins and dynamical evolution of these stars. Moreover, this supports the statement that the Eos Cloud is not a site of active star formation but rather a population of field stars similar to other regions at this Galactic latitude.
@@ 3.4 Potential for future star formation
Burkhart et al. (2025) calculated the overall Jeans stability of the Eos cloud, finding it to be marginally stable. They also calculated a global net molecular hydrogen dissociation and formation rate, finding that the cloud is net dissociating. Karoly et al. (2025) also concluded that the magnetic field strength is competitive with turbulence and gravity for the cloud. Overall these initial expectations are consistent with our conclusions drawn based on the Gaia populations towards the cloud. However, molecular clouds/star formation are heirarchical processes and so an analysis of the stability and dynamics, as well as the molecular formation and dissociation rate on a range of spatial scales (e.g. Krumholz et al., 2025) is required to fully conclude on the future star forming potential, or lack thereof, of the cloud. In particular, the presence of a CO-bright component (MBM 40) could be evidence of the beginning of collapse of a subset of the cloud. However, our study is in agreement with previous studies of the MBM 40 sub-region of the larger Eos cloud, indicating it is not currently star forming (Magnani et al., 1996), and Karoly et al. (2025) determined a magnetic field strength in MBM 40 which is also competitive with gravity and turbulence. Further detailed study of the gas conditions and kinematics are required to fully determine the future star formation potential of the cloud.
@@ 4 Summary and Conclusions
i) Comparison of the observed Gaia CMD with pre-main sequence model isochrones imply there are no significant populations less than tens of Myr old within 25 degrees of the centre of the Eos cloud, over a distance 70-150 pc from Earth (note that the Eos cloud extends from $\sim 95-136$ pc).

### v2

@@ 3.1 Color Magnitude Diagram
To analyze the age distribution of stars in the Eos Cloud, we constructed a Color-Magnitude Diagram (CMD) using Gaia DR3 photometric data. The CMD plots the absolute G magnitude ($M_{\textrm{G}}$) against the ($G_{\textrm{BP}}$-$G_{\textrm{RP}}$)0 color index, providing a clear visualization of the stellar population in the region. Figure 2 shows a CMD for our extinction, reddening and parallax-corrected stellar sample. Overlaid are Baraffe et al. (2015) PMS evolutionary tracks and isochrones. Other star forming regions on the edge of the local supernova-driven bubble have typical stellar ages $<3$ Myr (e.g. for Taurus, Lupus, Ophiuchus, Chameleon; Comerón, 2008; Galli et al., 2021; Miret-Roig et al., 2022; Luhman, 2025). Conversely, towards the Eos cloud there are no stars at all in that age range over the $70-150$ pc distance we considered (Eos spans 94-136 pc). There is a population of $<20$ Myr stars, however we will demonstrate in section 3.3 that a population similar to this is found in other similar search volumes at this galactic latitude. Furthermore, it is worth noting that the model isochrones are for single stars and that, at a given age, equal-mass binary systems will form a sequence parallel to the single star sequence but $\sim$0.75 mag brighter. The Gaia CMD hence demonstrates convincingly that star formation is not currently actively taking place in the Eos cloud.
@@ 3.3 Adjacent Regions
The absence of clustering effects and the uniformity across all the regions as mentioned above, suggest that possibly the stars in the Eos cloud and its surroundings have been dynamically dispersed over time. Possible mechanisms for this dispersion include Galactic tidal interactions, past dynamical mixing, or dissolution of an older stellar association (Bland-Hawthorn & Gerhard, 2016). Future studies incorporating full three-dimensional velocity data (e.g., using radial velocity measurements from Gaia DR4 or spectroscopic surveys) could provide deeper insights into the origins and dynamical evolution of these stars. Moreover, this supports the statement that the Eos Cloud is not a site of active star formation but rather contains a population of field stars similar to other regions at this Galactic latitude.
@@ 3.4 Potential for future star formation
Burkhart et al. (2025) calculated the overall Jeans stability of the Eos cloud, finding it to be marginally stable. They also calculated a global net molecular hydrogen dissociation and formation rate, finding that the cloud is net dissociating. Karoly et al. (2025) also concluded that the magnetic field strength is competitive with turbulence and gravity for the cloud. Overall these initial expectations are consistent with our conclusions drawn based on the Gaia populations towards the cloud. However, molecular clouds/star formation are hierarchical processes and so an analysis of the stability and dynamics, as well as the molecular formation and dissociation rate on a range of spatial scales (e.g. Krumholz et al., 2025) is required to fully conclude on the future star forming potential, or lack thereof, of the cloud. In particular, the presence of a CO-bright component (MBM 40) could be evidence of the beginning of collapse of a subset of the cloud. However, our study is in agreement with previous studies of the MBM 40 sub-region of the larger Eos cloud, indicating it is not currently star forming (Magnani et al., 1996), and Karoly et al. (2025) determined a magnetic field strength in MBM 40 which is also competitive with gravity and turbulence. Further detailed study of the gas conditions and kinematics are required to fully determine the future star formation potential of the cloud.
@@ 4 Summary and Conclusions
i) Comparison of the observed Gaia CMD with pre-main sequence model isochrones imply there are no significant populations less than tens of Myr old within 25 degrees of the centre of the Eos cloud, over a distance 70-150 pc from Earth (note that the Eos cloud extends from $\sim 94-136$ pc).
@@ Acknowledgements
We thank the referee for their timely and positive review of our manuscript.